A train of mass 4.0x10 kg runs south wards at a speed of 20m/s at a latitude 60 degree N what is the magnitude and direction of the coriolis force on tha train?


1
Expert's answer
2021-08-19T14:23:07-0400

Given:

"m=4*10\:\rm kg"

"v=20\:\rm m/s"

"\lambda=60^{\circ}"


The magnitude of the Coriolis force


"F_c=2mv\Omega\sin\lambda"

"F_c=2*4*10*20*\frac{2\pi}{24*3600}\sin 60^{\circ}=0.1\:\rm N"

The direction is northward.